A group of 10 people includes 3 retirees, 2 part-time workers, 2 discouraged workers, 1 unemployed worker, and 2 full-time workers.
If the discouraged workers in this group became active job searchers, the:

unemployment rate would fall.

labor force participation rate would fall.

unemployment rate and the labor force participation rate would rise.

unemployment rate and the labor force participation rates would both fall.
Answer #1

Answer

the adult population =total people - retirees=10-3=7

Employed =part-time workers+full-time workers=2+2=4

Unemployed =1

Labor force =employed + unemployed=4+1=5

Unemployent rate =(unemployed /labor force)*100

=(1/5)*100=20%

=======

after the discouraged workers in this group became active job searchers then

unemployed =1+2=3

labor force =4+3=7

Unemployment rate =(3/7)*100

=42.8571429

============

LFPR before the change =(LF/working age population)*100=(5/7)*100=71.4285714

after change =(7/7)*100=100%

=========

so in total

unemployment rate increases

LFPR increases

Option 3 is correct

unemployment rate and the labor force participation rate would rise.